OpenStudy (anonymous):

How can you tell what a metal is or its atomic weight just by knowing the specific heat.

OpenStudy (anonymous):

The specific heat is approximately \[C = 3Nk_B = 3 n R\] so you can find the number of moles in a sample: \[ n = \frac{C}{3R} \] If you know the mass of the sample (say, M) then the molar mass is just \[ \frac{M}{n} = \frac{3MR}{C} \]